OpenStudy (anonymous):

How do I expand and simplify the first 3 terms of: (3b^2 - 2/b)^14?

OpenStudy (anonymous):



OpenStudy (anonymous):

do you know the pascal coefficients for this?

OpenStudy (anonymous):

1 14 91 364 1001 2002 3003 3432 3003 2002 1001 364 91 14 1 \[(a+b)^{14}=a^{14}+14a^{13}b+91a^{12}b^2...\]

OpenStudy (anonymous):

Now just let a=3b^2 and b=-2/b :)

OpenStudy (anonymous):

Bad choice of letters on my part :/

OpenStudy (anonymous):

or x= 3b..?

OpenStudy (anonymous):

\[(x+y)^{14}=x^{14}+14x^{13}y+91x^{12}y^2...\] let x=3b^2 and y=-2/b :)